# In LIGO’s Sight? Vigorous Coherent Gravitational Waves from Cooled Collapsar Disks

### v1

@@ Abstract
We present the first numerical study of gravitational waves (GWs) from collapsar disks, using state-of-the-art 3D general-relativistic magnetohydrodynamic simulations of collapsing stars. These simulations incorporate a fixed Kerr metric for the central black hole (BH) and employ simplified prescriptions for disk cooling. We find that cooled disks with an expected scale height ratio of $H/R\gtrsim 0.1$ at $\sim 10$ gravitational radii induce Rossby instability in compact, high-density rings. The trapped Rossby vortices generate vigorous coherent emission regardless of disk magnetization and BH spin. For BH mass of $\sim 10\,M_{\odot}$, the GW spectrum peaks at $\sim 100\,{\rm Hz}$ with some breadth due to various non-axisymmetric modes. The spectrum shifts toward lower frequencies as the disk viscously spreads and the circularization radius of the infalling gas increases. Weaker-cooled disks with $H/R\gtrsim 0.3$ form a low-density extended structure of spiral arms, resulting in a broader, lower-amplitude spectrum. Assuming an optimistic detection threshold with a matched-filter SNR of 20 and a rate similar to supernovae Ib/c, LIGO-Virgo-KAGRA (LVK) could detect $\lesssim 1$ event annually, suggesting that GW events may already be hidden in observed data. Third-generation GW detectors could detect dozens to hundreds of collapsar disks annually, depending on the cooling strength and the disk formation rate. The GW amplitudes from collapsar disks are $\gtrsim 100$ times higher with a substantially greater event rate than those from core-collapse supernovae, making them potentially the most promising burst-type GW class for LVK and Cosmic Explorer. This highlights the importance of further exploration and modeling of disk-powered GWs, promising insights into collapsing star physics.
@@ 1 Introduction
Similar to inspiralling sources, burst-type sources are astrophysical events that generate transient GWs. However, their GW emission is less coherent, posing challenges for creating deterministic waveform models and, consequently, for their detection. A promising site for such burst-type sources is powerful explosions such as core-collapse supernovae (CCSNe). However, numerical simulations have shown that due to their quasi-spherical structure, only a minute fraction of the CCSN energy is released in GWs, where the dominant mechanism is stochastic excitation of the protoneutron star (PNS) f- and g-modes, which are strongly coupled to gravitational radiation (e.g., Nakamura et al., 2016; Andresen et al., 2017, 2019, 2021; Yakunin et al., 2017; Morozova et al., 2018; Radice et al., 2019; Abdikamalov et al., 2020; Mezzacappa et al., 2020, 2023; Vartanyan et al., 2023; Mezzacappa & Zanolin, 2024; Powell & Müller, 2024; Richardson et al., 2024). The predicted small excitation amplitudes imply that even the third-generation GW detectors can only detect GWs powered by PNS accretion in the Milky Way and satellite galaxies at design sensitivity (Srivastava et al., 2019; Evans et al., 2023; Corsi et al., 2024).
@@
If the progenitor star has sufficient angular momentum, its collapse will form an accretion disk around a newly formed black hole (BH, “collapsar”). Accretion disks have been considered promising GW emitters due to various instabilities (Kobayashi & Mészáros, 2003; van Putten & Levinson, 2003; Levinson et al., 2015; van Putten et al., 2019). If a global instability develops, non-axisymmetric modes could give rise to density waves, introducing a quadrupole moment that emits GWs. One criterion to examine the stability of the disk is by comparing its pressure and gravity. The pressure gradient can counteract the gravitational forces trying to pull material inward, thereby preventing excessive collapse or disruption. If the disk becomes gravitationally unstable (Chen & Beloborodov, 2007), it may fragment into high-density clumps (Gammie, 2001), which might give rise to strong GW emission (Piro & Pfahl, 2007). However, it remains to be seen whether such accretion disks exist in astrophysical systems and the extent of their asymmetry, highlighting the need for general-relativistic magnetohydrodynamic (GRMHD) simulations.
@@
Previous numerical attempts to assess the GW emission from accretion disks employed a preset self-gravitating torus setup (Kiuchi et al., 2011; Korobkin et al., 2011; Wessel et al., 2021). These studies found that non-axisymmetric modes emerge in hydrodynamic simulations due to Papaloizou-Pringle instabilities (PPI), which are quenched by magnetorotational instabilities (MRI) when magnetic fields are present (Bugli et al., 2018; Wessel et al., 2023). However, the relevance of torus studies to astrophysical systems such as collapsars is questionable for several reasons: (1) Ad-hoc prescribed tori evolve differently than accretion disks that form self-consistently during the collapse; (2) Torus simulations lack the constant gas supply onto the disk and the disk-envelope interaction; (3) Previous torus studies have not considered cooling effects, which are expected in the high-density collapsar disks due to neutrino emission (e.g., Narayan et al., 2001; van Putten & Levinson, 2003; Kohri et al., 2005; Chen & Beloborodov, 2007; Just et al., 2022). Cooling alters the disk thermodynamics and increases its density, potentially making it the primary driving mechanism for the disk to become gravitationally unstable (Batta & Lee, 2014), leading to vigorous GW emission.
@@
Another potentially important source of density perturbations is the Rossby wave instabilities (RWIs Lovelace et al., 1999; Tagger & Pellat, 1999; Li et al., 2000, 2001). RWIs emerge due to sharp gradients in the disk, which act as a potential well that can trap and amplify coherent structures like Rossby vortices. Therefore, RWIs are expected to be particularly strong in transiently forming disks. This insight motivates detailed first-principles simulations of self-consistently forming disks without considering self-gravity. Indeed, we find that the overdense Rossby vortices appear robustly in our simulations, creating a substantial mass quadrupole, responsible for generating GWs that may be detectable in LVK. The vigorous RWI-powered GWs were overlooked in previous studies because their setups lacked cooling mechanisms, necessary to increase the disk mass density and facilitate the formation of RWIs.
@@
In this letter, first in a series of papers, we begin to build the case for serious and intense searches for GW bursts from collapsar disks by current and future GW interferometer teams. The structure of the letter is as follows. We outline the numerical setup and the GW calculation methodology in §2. In §3, we describe the disk evolution and the emergence of instabilities that give rise to non-axisymmetric modes. Subsequently, in §4, we present the resultant vigorous GW signal, followed by a discussion on the expected event rates in LVK and third-generation detectors in §5. We deliberate on the implications of our results and conclude in §6.
@@ 2.1 Numerical setup
A robust GW signal may emerge from compact, dense accretion disks. Such conditions are prevalent in accretion disks around BHs formed due to the gravitational collapse of a rotating stellar core. Similar conditions also exist in disks formed following a merger of two compact objects, at least one of which is an NS. Both of these systems exhibit high mass accretion rates of $\dot{M}\gtrsim 10^{-2}\,\,{M_{\odot}}$ s-1, making them susceptible to neutrino cooling. Analytic models (Chen & Beloborodov, 2007) and numerical simulations incorporating neutrino transport (Siegel et al., 2019; Foucart, 2023) suggest that neutrino-cooled disks cool down to a characteristic ratio of disk height to radius of $0.1\lesssim H/R\lesssim 0.3$, depending on the effective viscosity and BH spin. $H/R$ tends to be smaller, $H/R\approx 0.1$, in the innermost disk regions, where, as we shall see, most of the GW radiation is emitted. Neutrino transport schemes are computationally expensive and thus do not allow following the disk evolution over long timescales. Therefore, our simulations do not incorporate full neutrino transport, but a cooling prescription that artificially maintains a fixed $H/R$ (Noble et al., 2009). This prescription mimics radiative cooling by reducing the internal energy of the disk over the local Keplerian timescale until it matches a target disk scale height.
@@
The fiducial collapsar simulations (model $C$ in table 1) adhere to the setup outlined in Gottlieb et al. (2022b), employing a stellar mass of $M_{\star}=20\,\,{M_{\odot}}$, and maximum initial magnetization ($\sigma_{0}\equiv b^{2}/4\pi\rho c^{2}$, where $b$ is the comoving magnetic field strength and $\rho$ is the comoving mass density) within the stellar core capped at ${\rm max}(\sigma_{0})=10^{-3}$. The angular momentum profile provides a fixed circularization radius of $r_{\rm circ}=25\,r_{g}$, similar to what is suggested by stellar evolution models during the first several seconds (Gottlieb et al. 2024, in prep.). The disk scale height is imposed to maintain $H/R=0.1$ at all times. The BH mass and dimensionless spin parameter are $\,{M_{\rm BH}}=10\,\,{M_{\odot}}$, and $\,{a_{\rm BH}}=0.8$, respectively. In addition to our fiducial model $C$, we also consider a slowly-spinning BH with $\,{a_{\rm BH}}=0.1$ (model $Ca$) and a weaker $B$-field (model $Cb$). Both exhibit effectively thicker disks which increase in height from $H/R=0.1$ close to the BH to $H/R\gtrsim 0.2$ at $R\gtrsim 20\,r_{g}$, consistent with neutrino-transport collapsar simulations (Issa et al. 2024, in prep.). These disks are thicker because less gravitational energy (due to increased ISCO in $Ca$) or magnetic dissipation ($Cb$) is available to heat, making the cooling process less efficient. We also consider a weakly-cooled disk with $H/R=0.3$ in model $Cc$ which increases to $H/R\approx 0.4$ at $r=20\,r_{g}$, similar to a collapsar disk without cooling in Gottlieb et al. (2022a). We find that GWs from weakly-cooled magnetized disks with $H/R=0.3$ are outshone by the cocoon-powered GWs. Therefore, to investigate the GW signature of the weakly-cooled disk, we impose $b=0$ in model $Cc$ such that no electromagnetically-driven jet and cocoon form. We outline the list of simulations in Table 1.
@@
and is imposed to be divergence-free. The potential normalization is chosen so the gas-to-magnetic pressure ratio is $\beta_{p}\approx 10^{4}$.
We conduct the simulations using the 3D GPU-accelerated code h-amr (Liska et al., 2022), employing an ideal equation of state with an adiabatic index of $4/3$, suitable for radiation-dominated gas. The simulations adopt spherical polar coordinates, $r$, $\theta$, $\phi$ with a logarithmic cell distribution along the $r$-direction, spanning from just inside the event horizon to $R_{\rm max}$, which for collapsars is located outside of the star radius (see Tab. 1). Along $\theta$- and $\phi$-directions, the cells are distributed uniformly. In total, the base grid comprises $N_{r}\times N_{\theta}\times N_{\phi}$ cells in the $r$-, $\theta$- and $\phi$-directions, respectively. For the collapsar and BNS merger simulations, $N_{r}=320,192$; $N_{\theta}=96,96$; and $N_{\phi}=192,96$, respectively. We incorporate local adaptive time-stepping and static mesh refinement. One (two) level of refinement is applied in the innermost $100\,(50)\,r_{g}$ in all collapsar (merger) simulations. We ensure that the wavelength of the fastest-growing MRI mode in the magnetohydrodynamic simulations is adequately resolved when the disk reaches a steady state evolution. In particular, we find that the MRI quality factor $Q_{\rm MRI}$, denoting the number of cells per the MRI wavelength, satisfies $Q_{\rm MRI}\gtrsim 30$ in both the $\hat{\theta}$ and $\hat{\phi}$ directions, surpassing the minimum requirement of $Q_{\rm MRI}\sim 10$ (e.g., Hawley et al., 2011).
@@ 2.2 GW approximations
where $\tilde{m}$ is the non-axisymmetric mode. Eq. (5) implies that the multipole expansion remains valid when $R_{d}\gtrsim\tilde{m}^{2}r_{g}$. For the quadrupole moment ($\tilde{m}=2$), this corresponds to $\,{f_{\rm GW}}\lesssim 800\,{\rm Hz}$.
@@ 2.2.2 Post-process calculation
and $\tilde{h}_{\times}(f)$ and $\tilde{h}_{+}(f)$ represent the strain polarizations in the frequency domain. To evaluate the signal detectability and event rate, we estimate the matched-filter signal-to-noise ratio (SNR) as (Flanagan & Hughes, 1998; Moore et al., 2015):
@@
where $S(f)$ denotes the noise power spectral density of the GW detector, for which we use the PyCBC package  (Usman et al., 2016; Nitz et al., 2021) for the noise power-spectral densities. $\tilde{h}_{A}(f)$ is the dimensionless GW strain, defined by:
@@ 3 Hydrodynamic evolution
Figure 1: 3D renderings illustrate the logarithmic mass density (in .c.g.s.) of collapsar accretion disks. The length scale can be measured with respect to the BH event horizon radius, represented by a black circle. (a) The early formation stages of the disk ($\sim 1\,\,{{\rm s}}$) feature low-density spiral density waves. (b) With the accumulation of infalling gas, a high-density ring forms at $R\approx 15\,r_{g}$. RWIs in the disk give rise to non-axisymmetric modes. Over time, the disk viscously spreads and its density progressively drops. (c) In the absence of strong cooling, the disk exhibits an extended structure [note that the horizon size in panels (a) and (b) is considerably larger than that in panel (c)], and attains lower densities compared to strongly-cooled disks. Consequently, weakly-cooled disks exhibit spiral arms at all times, giving rise to a weaker GW signal. Full movies showcasing the evolution of all collapsar accretion disks are available at http://www.oregottlieb.com/disk_gw.html.
The first stellar shell with a circularization radius exceeding the innermost stable circular orbit (ISCO) radius, will form an accretion disk. We simulate the collapse of a spherically symmetric mass distribution star, incorporating a solid body rotation profile that facilitates the rapid formation of an accretion disk. Figure 1 presents 3D renderings depicting the logarithmic mass density at different phases of strongly-cooled disk formation (panels a,b) and for weakly-cooled disks (c). The strongly-cooled accretion disks (models $C,Ca,Cb$) exhibit similar evolution. We identify two stages in the formation of strongly-cooled disks in collapsars. As illustrated in Figure 1(a), during the initial stages of disk formation ($\lesssim 1\,\,{{\rm s}}$), the disk is composed of low-density plasma. In low-density environments, pressure and density waves propagate freely, enabling differential rotation to amplify small perturbations into large, coherent spiral arms. Consequently, at the onset of disk formation, the radial expansion of the spiral waves results in a roughly flat and attenuated GW spectrum.
@@
Figure 2: (a) Velocity streamlines in the co-rotating frame of the high-density wave at $(x,y)\approx(-150,100)\,\,{{\rm km}}$ portray a trapped Rossby vortex. (b) The mass accretion rate on the BH horizon remains roughly constant in time and across models. (c) The jet efficiency on the BH horizon is substantial for rapidly spinning BHs with strong magnetic fields.
@@
Stellar evolution models suggest that the circularization radius remains roughly constant during the first several seconds at $R\sim 20\,r_{g}$ (Gottlieb et al. 2024, in prep.), as employed in our simulations. Nevertheless, the presence of strong magnetic fields facilitates MRI, which transports angular momentum outwards, thereby driving disk expansion and reducing disk density. This implies that later falling gas will merge with the outer parts of the disk before reaching its circularization radius. After several seconds, stellar evolution models (e.g., Woosley & Heger, 2006) suggest that the circulation radius of the gas increases as a function of the radius at the time of collapse, and thus the disk expansion will become faster. We leave a detailed exploration of the impact of the stellar rotational profile on the disk structure for future work.
@@
where the integration is calculated on the BH horizon $r_{h}$, and only for fluid elements with magnetization $\sigma>1$, $g$ is the metric determinant, $T^{r}_{t}$ denotes the radial energy flux density component of the mixed stress-energy tensor $T$, and $u^{\mu}$ is the four-velocity such that $\rho u^{r}$ represents the radial mass-energy flux density. Model $C$ exhibits high jet efficiency through the Blandford-Znajek mechanism (BZ; Blandford & Znajek, 1977). In contrast, slowly rotating and weakly-magnetized BHs can eject merely $\eta_{j}\lesssim 1\%$ of the accreted power due to a smaller angular momentum reservoir or weaker magnetic flux on the horizon. We note that the emergence of collapsar jets in weakly-magnetized disks is observed only when cooling is strong (see Gottlieb et al., 2022a, for the absence of jets in weakly-magnetized thick disks).

### v2

@@ Abstract
We present the first numerical study of gravitational waves (GWs) from collapsar disks, using state-of-the-art 3D general relativistic magnetohydrodynamic simulations of collapsing stars. These simulations incorporate a fixed Kerr metric for the central black hole (BH) and employ simplified prescriptions for disk cooling. We find that cooled disks with an expected scale height ratio of $H/R\gtrsim 0.1$ at $\sim 10$ gravitational radii induce Rossby instability in compact, high-density rings. The trapped Rossby vortices generate vigorous coherent emission regardless of disk magnetization and BH spin. For BH mass of $\sim 10\,M_{\odot}$, the GW spectrum peaks at $\sim 100\,{\rm Hz}$ with some breadth due to various nonaxisymmetric modes. The spectrum shifts toward lower frequencies as the disk viscously spreads and the circularization radius of the infalling gas increases. Weaker-cooled disks with $H/R\gtrsim 0.3$ form a low-density extended structure of spiral arms, resulting in a broader, lower-amplitude spectrum. Assuming an optimistic detection threshold with a matched-filter signal-to-noise ratio of 20 and a rate similar to Type Ib/c supernovae, LIGO–Virgo–KAGRA (LVK) could detect $\lesssim 1$ event annually, suggesting that GW events may already be hidden in observed data. Third-generation GW detectors could detect dozens to hundreds of collapsar disks annually, depending on the cooling strength and the disk formation rate. The GW amplitudes from collapsar disks are $\gtrsim 100$ times higher with a substantially greater event rate than those from core-collapse supernovae, making them potentially the most promising burst-type GW class for LVK and Cosmic Explorer. This highlights the importance of further exploration and modeling of disk-powered GWs, promising insights into collapsing star physics.
@@ 1 Introduction
Similar to inspiraling sources, burst-type sources are astrophysical events that generate transient GWs. However, their GW emission is less coherent, posing challenges for creating deterministic waveform models and, consequently, for their detection. A promising site for such burst-type sources is powerful explosions such as core-collapse supernovae (CCSNe). However, numerical simulations have shown that due to their quasi-spherical structure, only a minute fraction of the CCSN energy is released in GWs, where the dominant mechanism is stochastic excitation of the proto-neutron star (PNS) $f$- and $g$-modes, which are strongly coupled to gravitational radiation (e.g., Nakamura et al., 2016; Andresen et al., 2017, 2019, 2021; Yakunin et al., 2017; Morozova et al., 2018; Radice et al., 2019; Abdikamalov et al., 2020; Mezzacappa et al., 2020, 2023; Vartanyan et al., 2023; Mezzacappa & Zanolin, 2024; Powell & Müller, 2024; Richardson et al., 2024). The predicted small excitation amplitudes imply that even the third-generation GW detectors can only detect GWs powered by PNS accretion in the Milky Way and satellite galaxies at design sensitivity (Srivastava et al., 2019; Evans et al., 2023; Corsi et al., 2024).
@@
If the progenitor star has sufficient angular momentum, its collapse will form an accretion disk around a newly formed black hole (BH; “collapsar”). Accretion disks have been considered promising GW emitters due to various instabilities (Kobayashi & Mészáros, 2003; van Putten & Levinson, 2003; Levinson et al., 2015; van Putten et al., 2019). If a global instability develops, nonaxisymmetric modes could give rise to density waves, introducing a quadrupole moment that emits GWs. One criterion to examine the stability of the disk is by comparing its pressure and gravity. The pressure gradient can counteract the gravitational forces trying to pull material inward, thereby preventing excessive collapse or disruption. If the disk becomes gravitationally unstable (Chen & Beloborodov, 2007), it may fragment into high-density clumps (Gammie, 2001), which might give rise to strong GW emission (Piro & Pfahl, 2007; Siegel et al., 2022). However, it remains to be seen whether such accretion disks exist in astrophysical systems and the extent of their asymmetry, highlighting the need for general-relativistic magnetohydrodynamic (GRMHD) simulations.
@@
Previous numerical attempts to assess the GW emission from accretion disks employed a preset self-gravitating torus setup (Kiuchi et al., 2011; Korobkin et al., 2011; Wessel et al., 2021). These studies found that nonaxisymmetric modes emerge in hydrodynamic simulations owing to Papaloizou-Pringle instabilities, which are quenched by magnetorotational instabilities (MRI) when magnetic fields are present (Bugli et al., 2018; Wessel et al., 2023). However, the relevance of torus studies to astrophysical systems such as collapsars is questionable for several reasons: (1) Ad-hoc prescribed tori evolve differently than accretion disks that form self-consistently during the collapse; (2) Torus simulations lack the constant gas supply onto the disk and the disk-envelope interaction; (3) Previous torus studies have not considered cooling effects, which are expected in the high-density collapsar disks owing to neutrino emission (e.g., Narayan et al., 2001; van Putten & Levinson, 2003; Kohri et al., 2005; Chen & Beloborodov, 2007; Just et al., 2022). Cooling alters the disk thermodynamics and increases its density, potentially making it the primary driving mechanism for the disk to become gravitationally unstable (Batta & Lee, 2014), leading to vigorous GW emission.
@@
Another potentially important source of density perturbations is the Rossby wave instabilities (RWIs Lovelace et al., 1999; Tagger & Pellat, 1999; Li et al., 2000, 2001). RWIs emerge as a result of sharp gradients in the disk, which act as a potential well that can trap and amplify coherent structures like Rossby vortices. Therefore, RWIs are expected to be particularly strong in transiently forming disks. This insight motivates detailed first-principles simulations of self-consistently forming disks without considering self-gravity. Indeed, we find that the overdense Rossby vortices appear robustly in our simulations, creating a substantial mass quadrupole, responsible for generating GWs that may be detectable in LVK. The vigorous RWI-powered GWs were overlooked in previous studies because their setups lacked cooling mechanisms, necessary to increase the disk mass density and facilitate the formation of RWIs.
@@
In this letter, the first in a series of papers, we begin to build the case for serious and intense searches for GW bursts from collapsar disks by current and future GW interferometer teams. The structure of the letter is as follows. We outline the numerical setup and the GW calculation methodology in §2. In §3, we describe the disk evolution and the emergence of instabilities that give rise to nonaxisymmetric modes. Subsequently, in §4, we present the resultant vigorous GW signal, followed by a discussion on the expected event rates in LVK and third-generation detectors in §5. We deliberate on the implications of our results and conclude in §6.
@@ 2.1 Numerical setup
A robust GW signal may emerge from compact, dense accretion disks. Such conditions are prevalent in accretion disks around BHs formed owing to the gravitational collapse of a rotating stellar core. Similar conditions also exist in disks formed following a merger of two compact objects, at least one of which is a neutron star (NS). Both of these systems exhibit high mass accretion rates of $\dot{M}\gtrsim 10^{-2}\,\,{M_{\odot}}$ s-1, making them susceptible to neutrino cooling. Analytic models (Chen & Beloborodov, 2007) and numerical simulations incorporating neutrino transport (Siegel et al., 2019; Foucart, 2023) suggest that neutrino-cooled disks cool down to a characteristic ratio of disk height to radius of $0.1\lesssim H/Rx\lesssim 0.3$, depending on the effective viscosity and BH spin. $H/R$ tends to be smaller, $H/R\approx 0.1$, in the innermost disk regions, where, as we shall see, most of the GW radiation is emitted. Neutrino transport schemes are computationally expensive and thus do not allow following the disk evolution over long timescales. Therefore, our simulations do not incorporate full neutrino transport, but a cooling prescription that artificially maintains a fixed $H/R$ (Noble et al., 2009). This prescription mimics radiative cooling by reducing the internal energy of the disk over the local Keplerian timescale until it matches a target disk scale height.
@@
The fiducial collapsar simulations (model $C$ in table 1) adhere to the setup outlined in Gottlieb et al. (2022b), employing a stellar mass of $M_{\star}=20\,\,{M_{\odot}}$, and maximum initial magnetization ($\sigma_{0}\equiv b^{2}/4\pi\rho c^{2}$, where $b$ is the comoving magnetic field strength and $\rho$ is the comoving mass density) within the stellar core capped at ${\rm max}(\sigma_{0})=10^{-3}$. The angular momentum profile provides a fixed circularization radius of $r_{\rm circ}=25\,r_{g}$, similar to what is suggested by stellar evolution models during the first several seconds (Gottlieb et al., 2024). The disk scale height is imposed to maintain $H/R=0.1$ at all times. The BH mass and dimensionless spin parameter are $\,{M_{\rm BH}}=10\,\,{M_{\odot}}$, and $\,{a_{\rm BH}}=0.8$, respectively. In addition to our fiducial model $C$, we also consider a slowly-spinning BH with $\,{a_{\rm BH}}=0.1$ (model $Ca$) and a weaker $B$-field (model $Cb$). Both exhibit effectively thicker disks which increase in height from $H/R=0.1$ close to the BH to $H/R\gtrsim 0.2$ at $R\gtrsim 20\,r_{g}$, consistent with neutrino-transport collapsar simulations (Issa et al. 2024, in prep.). These disks are thicker because less gravitational energy (due to increased innermost stable circular orbit (ISCO) in $Ca$) or magnetic dissipation ($Cb$) is available to heat, making the cooling process less efficient. We also consider a weakly-cooled disk with $H/R=0.3$ in model $Cc$ which increases to $H/R\approx 0.4$ at $r=20\,r_{g}$, similar to a collapsar disk without cooling in Gottlieb et al. (2022a). We find that GWs from weakly-cooled magnetized disks with $H/R=0.3$ are outshone by the cocoon-powered GWs. Therefore, to investigate the GW signature of the weakly-cooled disk, we impose $b=0$ in model $Cc$ such that no electromagnetically-driven jet and cocoon form. We outline the list of simulations in Table 1.
@@
and is imposed to be divergence-free. The potential normalization is chosen so that the gas-to-magnetic pressure ratio is $\beta_{p}\approx 10^{4}$.
We conduct the simulations using the 3D GPU-accelerated code h-amr (Liska et al., 2022), employing an ideal equation of state with an adiabatic index of $4/3$, suitable for radiation-dominated gas. The simulations adopt spherical polar coordinates, $r$, $\theta$, $\phi$ with a logarithmic cell distribution along the $r$-direction, spanning from just inside the event horizon to $R_{\rm max}$, which for collapsars is located outside of the star radius (see Tab. 1). Along the $\theta$- and $\phi$-directions, the cells are distributed uniformly. In total, the base grid comprises $N_{r}\times N_{\theta}\times N_{\phi}$ cells in the $r$-, $\theta$- and $\phi$-directions, respectively. For the collapsar and binary NS (BNS) merger simulations, $N_{r}=320,192$; $N_{\theta}=96,96$; and $N_{\phi}=192,96$, respectively. We incorporate local adaptive time-stepping and static mesh refinement. One (two) level of refinement is applied in the innermost $100\,(50)\,r_{g}$ in all collapsar (merger) simulations. We ensure that the wavelength of the fastest-growing MRI mode in the magnetohydrodynamic simulations is adequately resolved when the disk reaches a steady state evolution. In particular, we find that the MRI quality factor $Q_{\rm MRI}$, denoting the number of cells per the MRI wavelength, satisfies $Q_{\rm MRI}\gtrsim 30$ in both the $\hat{\theta}$ and $\hat{\phi}$ directions, surpassing the minimum requirement of $Q_{\rm MRI}\sim 10$ (e.g., Hawley et al., 2011).
@@ 2.2 GW approximations
where $\tilde{m}$ is the nonaxisymmetric mode. Eq. (5) implies that the multipole expansion remains valid when $R_{d}\gtrsim\tilde{m}^{2}r_{g}$. For the quadrupole moment ($\tilde{m}=2$), this corresponds to $\,{f_{\rm GW}}\lesssim 800\,{\rm Hz}$.
@@ 2.2.2 Post-process calculation
with $\tilde{h}_{\times}(f)$ and $\tilde{h}_{+}(f)$ representing the strain polarizations in the frequency domain. To evaluate the signal detectability and event rate, we estimate the matched-filter signal-to-noise ratio (SNR) as (Flanagan & Hughes, 1998; Moore et al., 2015):
@@
where $S(f)$ denotes the noise power spectral density of the GW detector, for which we use the PyCBC package  (Usman et al., 2016; Nitz et al., 2021) for the noise power-spectral densities. Parameter $\tilde{h}_{A}(f)$ is the dimensionless GW strain, defined by:
@@ 3 Hydrodynamic evolution
Figure 1: 3D renderings illustrate the logarithmic mass density (in .c.g.s.) of collapsar accretion disks. The length scale can be measured with respect to the BH event horizon radius, represented by a black circle. (a) The early formation stages of the disk ($\approx 1\,\,{{\rm s}}$) feature low-density spiral density waves. (b) , with the accumulation of infalling gas, a high-density ring forms at $R\approx 15\,r_{g}$. RWIs in the disk give rise to nonaxisymmetric modes . Over time, the disk viscously spreads and its density progressively drops. (c) In the absence of strong cooling, the disk exhibits an extended structure [note that the horizon size in panels (a) and (b) is considerably larger than that in panel (c)], and attains lower densities compared to strongly-cooled disks. Consequently, weakly-cooled disks exhibit spiral arms at all times , giving rise to a weaker GW signal. Full movies showcasing the evolution of all collapsar accretion disks are available at http://www.oregottlieb.com/disk_gw.html.
The first stellar shell with a circularization radius exceeding the ISCO radius, will form an accretion disk. We simulate the collapse of a spherically symmetric mass distribution star, incorporating a solid body rotation profile that facilitates the rapid formation of an accretion disk. Figure 1 presents 3D renderings depicting the logarithmic mass density at different phases of strongly-cooled disk formation (panels a,b) and for weakly-cooled disks (c). The strongly-cooled accretion disks (models $C,Ca,Cb$) exhibit similar evolution. We identify two stages in the formation of strongly-cooled disks in collapsars. As illustrated in Figure 1(a), during the initial stages of disk formation ($\lesssim 1\,\,{{\rm s}}$), the disk is composed of low-density plasma. In low-density environments, pressure and density waves propagate freely, enabling differential rotation to amplify small perturbations into large, coherent spiral arms. Consequently, at the onset of disk formation, the radial expansion of the spiral waves results in a roughly flat and attenuated GW spectrum.
@@
Figure 2: (a) velocity streamlines in the corotating frame of the high-density wave at $(x,y)\approx(-150,100)\,\,{{\rm km}}$ portray a trapped Rossby vortex. (b) The mass accretion rate on the BH horizon remains roughly constant in time and across models. (c) The jet efficiency on the BH horizon is substantial for rapidly spinning BHs with strong magnetic fields.
@@
Stellar evolution models suggest that the circularization radius remains roughly constant during the first several seconds at $R\sim 20\,r_{g}$ (Gottlieb et al., 2024), as employed in our simulations. Nevertheless, strong magnetic fields facilitate MRI, which transports angular momentum outwards, thereby driving disk expansion and reducing disk density. This implies that later falling gas will merge with the outer parts of the disk before reaching its circularization radius. After several seconds, stellar evolution models (e.g., Woosley & Heger, 2006) suggest that the circulation radius of the gas increases as a function of the radius at the time of collapse, and thus the disk expansion will become faster. We leave a detailed exploration of the impact of the stellar rotational profile on the disk structure for future work.
@@
where the integration is calculated on the BH horizon $r_{h}$, and only for fluid elements with magnetization $\sigma>1$, $g$ is the metric determinant, $T^{r}_{t}$ denotes the radial energy flux density component of the mixed stress-energy tensor $T$, and $u^{\mu}$ is the four-velocity such that $\rho u^{r}$ represents the radial mass-energy flux density. Model $C$ exhibits high jet efficiency through the Blandford-Znajek mechanism (BZ; Blandford & Znajek, 1977). In contrast, slowly rotating and weakly-magnetized BHs can eject merely $\eta_{j}\lesssim 1\%$ of the accreted power owing to a smaller angular momentum reservoir or weaker magnetic flux on the horizon. We note that the emergence of collapsar jets in weakly-magnetized disks is observed only when cooling is strong (see Gottlieb et al., 2022a, for the absence of jets in weakly-magnetized thick disks).